OpenStudy (anonymous):

the derivative of f(x)=tan^-1(x)^3

OpenStudy (anonymous):

Ok d/(du) of tan-1(u) = 1/(1+u^2) (tan-1(x))^3 u = tan-1(x) du = 1/(1+x^2) derivative: 3(tan-1(x))^2 (1/(1+x^2) (don't forget chain rule - the du at the end)

OpenStudy (anonymous):

im still getting the wrong answer